Claire Gerber wants to buy 400 shares of Google, which is selling in the market for $534.14 a share. Rather than liquidate all her savings, she decides to borrow through her broker at 5 percent a year. Assume that the margin requirement on common stock is 50 percent. If the stock rises to $625 a share over the next year, calculate the dollar profit and percentage return that Claire would earn if she makes the investment with 50 percent margin. Contrast these figures to what she'd make if she uses no margin. Calculate the dollar net profit. Round the answers to the nearest dollar. Without Margin and With 50% Margin Calculate the return on investment. Round the answers to two decimal places. Without Margin and With 50% Margin
Answer #1

Without margin:

Dollar gain=400*(625-534.14)=36344

Return on Investment=36344/(400*534.14)=17.01%

With margin:

Margin=400*534.14*50%=106828

Own Investment=400*534.14*50%=1068280

Interest to be paid=106828*5%=5341.40

Dollar gain=400*(625-534.14)-5341.4=31002.60

Return on Investment=31002.6/106828=29.02%